SLITEX(1) SLITEX(1) NAME slitex−makeLaTeXslides SYNOPSIS slitex[firstline] DESCRIPTION Theslitexprogramisavariantoflatex(1),designedforgeneratingslides.Itisdescribedin AppendixAofLaTeX−ADocumentPreparationSystembyLeslieLamport,publishedbyAddison-Wesley.Allthecapabilitiesdescribedintex(1)arepresent. SEEALSO tex(1),latex(1) 7Jan92 1